Question: A card is drawn from a standard deck. Find each of the following conditional probabilities, use the formulas and show your work.


a) The card is a queen, given it is a face card.

b) The card is a king, given that it is a spade.

c) The card is a spade, given that it is a king.



1
Expert's answer
2022-03-23T17:03:51-0400

a) "P=\frac{4}{12}=\frac{1}{3}."

b) "P=\frac{1}{13}."

c) "P=\frac{1}{4}."